Discharging RC circuit. • In the RC circuit shown, the battery has fully charged the capacitor, so Q0 = C E. Then at t = 0 the switch is thrown from position a to b. The battery emf is 20.0 V, and the capacitance C = 1.02 μF. The current I is observed to decrease to 0.50 of its initial value in 40 μs. (a) What is the value of Q, the charge on the capacitor, at t = 0? (b) What is the value of R?  (c) What is Q at t = 60 μs? 



1
Expert's answer
2021-01-28T17:10:33-0500


a) At "t=0" the value of the charge on the capacitor will be:


"Q=CE=1.02\cdot10^{-6}\ F\cdot20\ V=20.4\cdot10^{-6}\ C=20.4\ \mu C."

b) Let's write the equation for current:


"I(t)=I_0e^{-\dfrac{t}{RC}}."

Since, "I=0.5I_0", we get:


"0.5I_0=I_0e^{-\dfrac{t}{RC}},""ln(0.5)=ln(e^{-\dfrac{t}{RC}}),""-\dfrac{t}{RC}=ln(0.5)=-ln2,""R=\dfrac{t}{Cln2}=\dfrac{40\cdot10^{-6}\ s}{1.02\cdot10^{-6}\ F\cdot ln2}=57\ \Omega."

c) Let's first find the time constant:


"\tau=RC=57\ \Omega\cdot1.02\cdot10^{-6}=58\cdot10{-6}\ s=58\ \mu s."

Then, we can find "Q" at "t=60\ \mu s":


"Q=Q_0e^{-\dfrac{t}{\tau}}=20.4\cdot10^{-6}\ C\cdot e^{-\dfrac{60\ \mu s}{58\ \mu s}}=7.2\ \mu C."